# Explainable Anomaly Detection in Images and Videos: A Survey

### v2

@@ III-A Explainable Image Anomaly Detection
With the recent rise of foundation model architectures, there exist several very recent works in image anomaly detection literature leveraging the extraordinary reasoning ability of large foundation models trained with massive amounts of data. There are a series of work using CLIP [77] (CLIPAD [51][52] AnomalyCLIP [52]; WinCLIP [50]; CLIP-SAM [78]) for few-shot/zero-shot image anomaly detection and localization. More recently, Segment Any Anomaly [53] introduces the ”Segment Any Anomaly + (SAA+)” framework, a novel approach for anomaly segmentation under the zero-shot setting that utilizes hybrid prompt regularization to improve the adaptability of foundational models. Instead of relying on domain-specific fine-tuning like traditional anomaly segmentation models, SAA+ leverages the zero-shot generalization capabilities of foundational models, assembling them to utilize diverse multi-modal prior knowledge for anomaly localization, and further adapts them using hybrid prompts derived from domain expert knowledge and target image context. The method’s explainability stems from its innovative use of hybrid prompts, which are derived from both domain expert knowledge and the specific context of the target image. The basic idea is to utilize the alignment property and extraordinary representation ability of text and visual features of CLIP for generic abnormality learning and discrimination, which makes the methods naturally explainable in the corresponding text prompt space. Nowadays, with the rapid development of LLMs and Large MultiModal Models (LMMs) [79], the outstanding reasoning ability of pretrained language models has been employed for enhanced anomaly detection and explanations. AnomalyGPT [54] leverages LVLM for Industrial Anomaly Detection by simulating anomalous images with textual descriptions and fine-tuning LVLMs with prompt embeddings for direct anomaly presence and location assessment without manual thresholds. This approach is explainable as it supports multi-turn dialogues, allowing users to engage with the model for detailed anomaly localization and understanding, facilitated by a lightweight decoder for fine-grained semantic capture and prompt learner integration for enhanced LVLM alignment. Myriad [55] introduces a novel large multimodal model, leveraging vision experts for enhanced industrial anomaly detection (IAD), which includes an expert perception module for embedding prior knowledge from vision experts and an expert-driven vision-language extraction module for domain-specific vision-language representation. This method enhances explainability by incorporating expert-generated anomaly maps into the learning process, thereby providing detailed descriptions of anomalies, which improves both detection accuracy and the comprehensibility of the detection process; [56] tests the GPT-4V [79] capabilities in generic anomaly detection via examples, and [57] adapts GPT-4V for zero-shot anomaly detection in images through granular region division, prompt designing, and text to segmentation, enhancing visual question answering for direct anomaly evaluation without prior examples. Its explainability stems from leveraging VQA to interpret and segment anomalies based on textual prompts and image analysis, providing intuitive and detailed anomaly identification and localization capabilities. All these existing LLM/LMM-based IAD methods are in the form of concatenation of the extracted text and image embeddings sent to LLM, as depicted in Fig. 2 . Given a question related to the anomaly, the LLM can determine whether the image contains any anomaly, where it is, detailed descriptions, and even explanations. Owing to the massive knowledge the LLM is trained on, it can naturally generate meaningful textual explanations and further explanations can be acquired by changing or adding prompts.
@@ III-B Explainable Video Anomaly Detection
Attention-based methods
@@
The key idea of attention-based methods is to capture the gradient. How to calculate the gradient and how to use them to explain the model are the novel parts of this kind of method. Pang et al. [61] used a simple approach, class activation mapping (CAM) [80], to calculate attention weights. They then localized the identified anomaly by looking at high activation values on the attention map. Purwanto et al. [62] used a self-attention mechanism to model the relationships of the nodes in spatio-temporal graphs. They construct conditional random fields as nodes and learn self-attention on them. Visualization of the heat map on the self-attention random fields reveals the location of the anomaly. Other papers, such as [23], also use attention to localize the anomaly and give an explanation based on that. Although the attention map indicates the importance of a specific region or features that play a significant role in the prediction, it cannot explain why a model makes a correct or wrong prediction. For example, in Fig. 3, the attention map indicates a larger region than the anomaly, which reveals little information. Also, as [81] discussed, a random ResNet can provide an attention map that can capture the foreground object. In this case, it cannot tell anything even if it gives an accurate prediction.
Reasoning-based methods
@@
Different from the attention-based method, reasoning-based methods focus on giving high-level explanations by providing some context reasoning techniques, such as knowledge graph, scene graph, and causal inference. They explain the model in a semantic way that considers the foreground objects and the background information together. As shown in Fig. 4, the frame information can be formulated as graphs, including spatial-temporal graph, knowledge graph, scene graph, and so on. The information provided by graph can not only benefit anomaly detection but also bring interpretability to models.
For example, [65] proposed a framework with a global monitor and a local monitor. The global monitor constructs a scene graph that captures the global information, and the local monitor focuses on the human pose. Each interaction is observed by calculating its semantic distance from a collection of nominal embeddings derived from the training data. Typically, the (subject, predicate, object) triplet is mapped to numerical vectors using a semantic embedding network. The goal of metric learning is to extract information regarding the absence or presence of anomalies from the interaction embeddings. During training, the positive instance is randomly selected from the nominal training set, and the negative instance is randomly sampled from an artificially generated set of anomalous interaction triplets, e.g., a person hits a person. The objective is to minimize the distance between the positive samples and to maximize the negative samples. Thus, during testing, when an anomaly occurs, it can give the true reason for the anomaly and output an interpretable result. Similarly, [63] also relied on scene graphs to interpret the anomaly. However, it used multiclass SVM and naive Bayesian for anomaly detection.
@@
Intrinsically interpretable methods
@@
where $\phi(\cdot)$ shows the heat map of each frame, and $f(\cdot)$ denotes the prediction frame and $x$ is the ground truth frame, and $\theta(\cdot)$ represents the projection of difference of prediction and groundtruth. For example, [14] trained a neural network for predicting future frames. The predicted frame is compared to the ground truth frame and results in a pixel-level error map. Notably, although the error map looks like the saliency map or attention map, they are essentially different. The attention map reveals the weight of each pixel, while the error map indicates the prediction error based on each pixel. Based on the error map, it is easy to locate the error region. The object classification and action recognition model are then applied to classify and explain the anomalies.
Figure 6: Visualization of the attributes that are applied in existing methods. Those attributes are object-level. Class indicates the object category, like person, bag; features indicate the feature of an object, such as white, young and size refers to the size of the bounding box.
On the other hand, attribute interpretation methods make use of object attributes to determine whether an object is normal or abnormal. As illustrated in Fig. 6, most of them adopt attributes that are based on prior knowledge, such as velocity, pose, motion, action, and location. For example, [70] proposed a framework that extracts a high-level depiction of each video segment utilizing networks for appearance and motion. They choose the motion angle and velocity as the attributes to help explain the model. The method for selecting exemplars is applied to choose a representative subset of video segments for a specified spatial region, whose object and its corresponding angle and velocity are kept in the memory bank. During testing, a video sequence is classified as an anomaly if it is different from exemplars. Thus, the explanation can be made by the high-level representation. In particular, the appearance and motion parts of the combined feature vector are mapped to high-level attributes using the last fully connected layer of their respective networks. There are some other papers using different attributes. [67] considered object class and corresponding features and actions. They judge anomaly by predicting anomaly scores of each visual concept. The reason the anomaly can be inferred from these concepts is if one concept is largely deviating from normal patterns. [69] regard velocity, pose, and deep as attributes to interpret anomaly. Similarly, these features provide predictions and explanations for anomalies. [68] paid more attention to interactions, so they encoded object category, interaction, and size of bounding box and trained a Gaussian Mixture Model (GMM) from these representations.
@@
IV Differences and Commons
IV-A Explainable methods that can only be exclusively applied to images
From a feasibility perspective, the image anomaly detection methods can be applied to video data as well due to the fact that a video is a sequence of images in essence. Nevertheless, considering the intrinsic characteristics of the video anomaly detection task, many of them can not work reasonably on the benchmark video anomaly detection datasets.
The input-perturbation-based image anomaly detection approaches that utilized pretrained classification models [37, 39, 38] can not work on video anomaly detection tasks. They can be applied to the tasks of video anomaly detection in a frame-wise fashion, but image classification dataset-pretrained networks fail to differentiate between normal and abnormal frames in that the abnormality in video anomaly detection lies in detailed frame patch abruptness or intra-frame variation. These can not be represented by pre-trained image classification models which concentrate more on the categorical information. There is a line of video anomaly detection works that perform video detection of anomalies at the level of individual objects, and it seems that the input-perturbation-based image anomaly detection approaches can be incorporated into them. Nevertheless, such techniques fail to detect anomalies that result from intra-frame variation, which occupies a large proportion of all the anomalies, including people running, skating, falling, fighting, etc. This is because the essence of such kind of anomaly is the relative alternation across the frames, not the object category difference.
For image OOD detection methods using generative models including [43, 44], they learn the density function of the in-distribution data so that the likelihood-related score will be lower for outliers. Such methods certainly can not work for video anomaly detection owing to the fact that they can neither distinguish between unseen object category and seen object category nor model the variation across frames.
In summary, an explainable image anomaly detection method can be adapted for videos only if they have an error-based anomaly score. The video data can be concatenated in channel dimension to feed into the model and the error-based score can still detect anomalies in video frames, either object anomalies or motion anomalies.
IV-B Explainable methods that can only be exclusively applied to videos
Video anomaly detection has primarily focused on detecting anomalies in actions. As a result, video-explainable methods have emerged to provide explanations by leveraging video-specific features, such as temporal information and action information. These model-specific explainable methods are tailored to videos, as images lack the contextual cues necessary for such analysis. Conversely, reasoning-based explainable methods heavily rely on the contextual information inherent in videos. For instance, scene graphs have been employed as a general tool to analyze object relationships. However, image anomaly detection suffers from the inherent limitation of having only a few objects present, thereby limiting the scene graph’s capacity to provide meaningful explanations for predictions due to the lack of contextual information.
In order to address this limitation, [65] have proposed a novel approach that constructs a scene graph capable of capturing global information. The authors utilize a semantic embedding network to map (subject, predicate, object) triplets to numerical vectors. Through metric learning, the aim is to extract information pertaining to the absence or presence of anomalies from the interaction embeddings. During testing, when an anomaly occurs, this approach provides the ability to identify the true cause of the anomaly and produce an interpretable result.
Another key aspect specific to videos is the temporal sequence. Many video anomaly detection methods [9, 83, 6] make use of both the RGB frame cube and the optical flow cube. The optical flow cube provides crucial temporal and action context that is not available in individual images. By leveraging this information, several explainable methods have explored temporal sequences to generate interpretable results.
In summary, video anomaly detection research has witnessed advancements in explainability techniques. While model-specific explainable methods rely on video-specific features, such as temporal and action information, reasoning-based explainable methods require contextual cues present in videos. The construction of scene graphs capturing global information and the exploitation of temporal sequences through optical flow analysis have further enhanced the interpretability of video anomaly detection approaches.
IV-C Commmons

### v4

@@ III-A4 Foundation-model-based IAD
With the recent rise of foundation model architectures, there exist several very recent works in image anomaly detection literature leveraging the extraordinary reasoning ability of large foundation models trained with massive amounts of data. A series of work has explored adapting CLIP [108] for few-shot or zero-shot image anomaly detection and localization, leveraging its aligned vision-language representations. Examples include CLIPAD [66], AnomalyCLIP [67], WinCLIP [65], and CLIP-SAM [73]. Recent advancements focus on enhancing zero-shot capabilities through object-agnostic prompt learning [74], anomaly-aware adaptations of CLIP [75], and Bayesian approaches to prompt learning [78]. The basic idea is to utilize the alignment property and extraordinary representation ability of text and visual features of CLIP for generic abnormality learning and discrimination, which makes the methods naturally explainable in the corresponding text prompt space. Some approaches incorporate learnable class names to aid detection, particularly in long-tailed scenarios [109]. Further, Segment Any Anomaly [68] introduces the ”Segment Any Anomaly + (SAA+)” framework, a novel approach for anomaly segmentation under the zero-shot setting that utilizes hybrid prompt regularization to improve the adaptability of foundational models. Instead of relying on domain-specific fine-tuning like traditional anomaly segmentation models, SAA+ leverages the zero-shot generalization capabilities of foundational models, assembling them to utilize diverse multi-modal prior knowledge for anomaly localization, and further adapts them using hybrid prompts derived from domain expert knowledge and target image context. The method’s explainability stems from its innovative use of hybrid prompts, which are derived from both domain expert knowledge and the specific context of the target image. Specific adaptations have also been developed for generalizable anomaly detection in medical images using VLMs [76].
Nowadays, with the rapid development of LLMs and Large MultiModal Models (LMMs) [110], the outstanding reasoning ability of pretrained language models has been employed for enhanced anomaly detection and explanations. AnomalyGPT [69] leverages LVLM for Industrial Anomaly Detection by simulating anomalous images with textual descriptions and fine-tuning LVLMs with prompt embeddings for direct anomaly presence and location assessment without manual thresholds. This approach is explainable as it supports multi-turn dialogues, allowing users to engage with the model for detailed anomaly localization and understanding, facilitated by a lightweight decoder for fine-grained semantic capture and prompt learner integration for enhanced LVLM alignment. Myriad [70] introduces a novel large multimodal model, leveraging vision experts for enhanced industrial anomaly detection (IAD), which includes an expert perception module for embedding prior knowledge from vision experts and an expert-driven vision-language extraction module for domain-specific vision-language representation. This method enhances explainability by incorporating expert-generated anomaly maps into the learning process, thereby providing detailed descriptions of anomalies, which improves both detection accuracy and the comprehensibility of the detection process; [71] tests the GPT-4V [110] capabilities in generic anomaly detection via examples, and [72] adapts GPT-4V for zero-shot anomaly detection in images through granular region division, prompt designing, and text to segmentation, enhancing visual question answering for direct anomaly evaluation without prior examples. Its explainability stems from leveraging VQA to interpret and segment anomalies based on textual prompts and image analysis, providing intuitive and detailed anomaly identification and localization capabilities. More recent works continue this trend, exploring training-free approaches [79], employing VLM-based text features for logical explanation [111], and leveraging MLLMs for explicit zero-shot reasoning about anomalies [81]. All these existing LLM/LMM-based IAD methods are typically structured by concatenating extracted text and image embeddings and feeding them into an LLM, as depicted in Fig. 2. Given a question related to the anomaly, the LLM can determine whether the image contains any anomaly, where it is, detailed descriptions, and even explanations. Owing to the massive knowledge the LLM is trained on, it can naturally generate meaningful textual explanations, and further clarifications can be acquired by changing or adding prompts. The integration of few-shot learning techniques with residual learning also shows promise for building generalist anomaly detection systems with potential explainability benefits [77], as does the development of unified models for few-shot detection [112].
@@ III-B Explainable Video Anomaly Detection
III-B1 Attention-based methods
@@
The key idea of attention-based methods is to capture the gradient. How to calculate the gradient and how to use them to explain the model are the novel parts of this kind of method. [85] used a simple approach, class activation mapping (CAM) [113], to calculate attention weights. They then localized the identified anomaly by looking at high activation values on the attention map. [86] used a self-attention mechanism to model the relationships of the nodes in spatio-temporal graphs. They construct conditional random fields as nodes and learn self-attention on them. Visualization of the heat map on the self-attention random fields reveals the location of the anomaly. Other papers, such as [37], also use attention to localize the anomaly and give an explanation based on that. Although the attention map indicates the importance of a specific region or features that play a significant role in the prediction, it cannot explain why a model makes a correct or wrong prediction. For example, in Fig. 3, the attention map indicates a larger region than the anomaly, which reveals little information. Also, as [114] discussed, a random ResNet can provide an attention map that can capture the foreground object. In this case, it can tell nothing even if it gives an accurate prediction.
III-B2 Reasoning-based methods
@@
Logic-based methods
Different from the attention-based method, logic-based methods focus on giving high-level explanations by providing some context reasoning techniques, such as knowledge graph, scene graph, and causal inference. They explain the model in a semantic way that considers the foreground objects and the background information together. As shown in Fig. 4, the frame information can be formulated as graphs, including spatial-temporal graph, knowledge graph, scene graph, and so on. The information provided by graph can not only benefit anomaly detection but also bring interpretability to models.
For example, [89] proposed a framework with a global monitor and a local monitor. The global monitor constructs a scene graph that captures the global information, and the local monitor focuses on the human pose. Each interaction is observed by calculating its semantic distance from a collection of nominal embeddings derived from the training data. Typically, the (subject, predicate, object) triplet is mapped to numerical vectors using a semantic embedding network. The goal of metric learning is to extract information regarding the absence or presence of anomalies from the interaction embeddings. During training, the positive instance is randomly selected from the nominal training set, and the negative instance is randomly sampled from an artificially generated set of anomalous interaction triplets, e.g., a person hits a person. The objective is to minimize the distance between the positive samples and to maximize the negative samples. Thus, during testing, when an anomaly occurs, it can give the true reason for the anomaly and output an interpretable result. Similarly, [87] also relied on scene graphs to interpret the anomaly. However, it took advantage of multiclass SVM and naive Bayesian for anomaly detection.
@@
LLM-based methods
LLM-based methods improve explainable video anomaly detection via integrating visual and linguistic information. VERA [91] employs learned reflective questions to guide vision-language models in generating anomaly scores, ensuring explainability through these questions. AnomalyRuler [92] detects anomalies through inducing rules from normal samples, with the rules themselves providing clear explanations. LAVAD [93] generates textual descriptions and leverages LLMs for reasoning, offering explanations via the descriptions and prompting process. Holmes-VAU [94] targets anomaly-rich regions using a multimodal LLM, delivering explanations at varying levels of granularity. VadCLIP aligns visual and linguistic features for precise detection, linking anomalies to descriptive language for explainability. Anomize [95] guides detection with textual information and label relations, explaining novel anomalies through these connections. [90] produces textual explanations using video-based LLMs, making the detection process transparent. These methods collectively offer diverse, explainable approaches to identifying anomalies in video data harnessing the strong reasoning ability of the LLM backbones which were pre-trained on a large amount of textual data.
III-B3 Intrinsically interpretable methods
@@
where $\phi(\cdot)$ shows the heat map of each frame, and $f(\cdot)$ denotes the prediction frame and $x$ is the ground truth frame, and $\theta(\cdot)$ represents the projection of difference of prediction and groundtruth. For example, [21] trained a neural network for predicting future frames. The predicted frame is compared to the ground truth frame and results in a pixel-level error map. Notably, although the error map looks like the saliency map or attention map, they are essentially different. The attention map reveals the weight of each pixel, while the error map indicates the prediction error based on each pixel. Based on the error map, it is easy to locate the error region. The object classification and action recognition models are applied to classify and explain the anomalies.
On the other hand, attribute interpretation methods make use of object attributes to determine whether an object is normal or abnormal. For example, [99] proposed a framework that extracts a high-level depiction of each video segment utilizing networks for appearance and motion. They choose the motion angle and velocity as the attributes to help explain the model. The method for selecting exemplars is applied to choose a representative subset of video segments for a specified spatial region, whose object and its corresponding angle and velocity are kept in the memory bank. During testing, a video sequence is classified as an anomaly if it is different from exemplars. Thus, the explanation can be made by the high-level representation. In particular, the appearance and motion parts of the combined feature vector are mapped to high-level attributes using the last fully connected layer of their respective networks. There are some other papers using different attributes. [96] considered object class and corresponding features and actions. They judge anomaly by predicting anomaly scores of each visual concept. The reason the anomaly can be inferred from these concepts is if one concept is largely deviating from normal patterns. [98] regard velocity, pose, and deep as attributes to interpret anomaly. Similarly, these features provide predictions and explanations for anomalies. [97] paid more attention to interactions, so they encoded object category, interaction, and size of bounding box and trained a Gaussian Mixture Model (GMM) from these representations.
@@
III-B4 Memory-based methods
Another relevant category, often overlapping with reconstruction or prediction-based approaches, is memory-based methods [100, 101]. These methods explicitly store representations of normal patterns (e.g., feature vectors, prototypes) in a memory module during training. During inference, test samples are compared against the stored memory items. Anomalies are detected based on their dissimilarity to the closest memory items or their inability to be well-reconstructed using a combination of memory items. Explainability can arise from identifying which memory items are closest to a normal sample versus an anomalous one, or visualizing the reconstruction process using memory items. For instance, if an anomalous sample cannot be well represented by any stored normal pattern, this provides an implicit explanation for its abnormality.
III-C Other Related Methods
Beyond the main categories discussed, several other lines of research contribute to or overlap with explainable visual anomaly detection, although explainability might not be their primary focus. These include feature-based methods, self-supervised approaches, Shapley values, rule-based explanations, and novel class discovery.
Feature-based Methods
Many modern anomaly detection methods operate in the feature space of pre-trained networks (often trained on large datasets like ImageNet [116]). Anomalies are detected based on deviations in these feature representations. While not always explicitly designed for explainability, the features themselves can offer some level of interpretation. These methods often fall into categories like:
  • Feature Reconstruction These methods train models (like autoencoders) to reconstruct normal features. Anomalies are detected by high reconstruction errors in the feature space [117].
  • Feature Distillation (Student-Teacher) A ”teacher” network (pre-trained on normal data or a larger dataset) guides a ”student” network. Anomalies cause discrepancies between student and teacher features [118, 119].
  • Feature Comparison/Matching These methods compare features of test samples to a memory bank of normal features or learn discriminative boundaries in the feature space [3].
  • Feature Discrimination Approaches like SimpleNet [120] use simple architectures on pre-trained features to discriminate between normal and anomalous patterns effectively.
Explainability in these methods often comes from analyzing which features contribute most to the anomaly score or visualizing the differences between anomalous features and the distribution of normal features. However, relating deep features directly back to interpretable concepts in the input space remains a challenge.
Self-Supervised Learning (SSL) Methods
Self-Supervised Learning (SSL) has become a dominant paradigm in anomaly detection. Methods like CutPaste [121] and those using synthetic anomalies [122] learn representations by solving pretext tasks on normal data. The assumption is that models trained on normal data transformations will struggle to generalize to anomalous data. Explainability here is indirect; the model’s failure on the pretext task for an anomalous input serves as the explanation. However, understanding why a specific anomaly causes the model to fail on the pretext task requires further investigation, as discussed in Sec. VII.
Shapley Value Methods
Originating from cooperative game theory, Shapley values offer a principled way to attribute the prediction outcome to individual input features (e.g., pixels or superpixels). They have been applied to explain anomalies detected by models like autoencoders [123] and PCA [124], often using approximations like SHAP (SHapley Additive exPlanations) [125, 126]. These methods provide pixel-level importance scores, offering fine-grained explanations but can be computationally expensive.
Rule-Based Explanations
Another direction aims to interpret unsupervised anomaly detection models by extracting explicit, rule-based explanations [127]. This involves dissecting the internal workings of potentially black-box models to generate human-understandable rules that characterize why certain inputs are flagged as anomalous.
Novel Anomaly Class Discovery
Explainability is also crucial in scenarios where the goal is not just to detect anomalies but also to discover and characterize potentially novel classes of anomalies, particularly in industrial settings [128]. Understanding the characteristics of newly discovered anomaly types is essential for process improvement and adaptation.
These related areas highlight the diverse strategies employed in visual anomaly detection and the varying degrees and types of explainability they offer.
IV Differences and Commonalities
IV-A Explainable methods predominantly applied to images
From a feasibility perspective, many image anomaly detection methods can be technically applied to video data frame-by-frame. However, considering the intrinsic characteristics of video anomalies (often involving motion and temporal context), several IAD approaches are less suitable or directly applicable to standard VAD benchmarks.
Input-perturbation-based IAD approaches designed for OOD detection or semantic anomaly detection using pretrained image classification models [51, 53, 52] struggle with VAD. While they can operate frame-wise, image classification networks are typically insensitive to the subtle temporal variations or motion patterns that define many video anomalies (e.g., unusual gaits, interactions). Applying them to object tracks within videos might capture object-level anomalies but would miss motion-based or event-based anomalies, which rely on inter-frame relationships. The explanations derived from input perturbations (sensitivity of classification score to input changes) are less meaningful when the underlying classifier cannot distinguish normal vs. abnormal temporal patterns.
Similarly, certain generative model approaches focused on static image likelihoods, like those learning density functions for typical image classification datasets [57, 58], are ill-suited for VAD. They primarily capture the distribution of static appearances and fail to model temporal dynamics or complex event structures. Explanations based on likelihood scores from such models would not capture the essence of temporal anomalies.
In summary, explainable IAD methods heavily reliant on static image features or pre-trained image classifiers often lack the necessary mechanisms to model and explain temporal dynamics crucial for VAD. Adapting them effectively requires significant modifications to incorporate temporal context, moving beyond simple frame-wise application.
IV-B Explainable methods predominantly applied to videos
Video anomaly detection often focuses on unusual actions or events unfolding over time. Consequently, explainable VAD methods frequently leverage video-specific cues like temporal context, motion information (e.g., optical flow), and object interactions, making them less directly applicable to static images.
Model-specific explanations derived from architectures designed for temporal modeling (e.g., RNNs, 3D CNNs, Transformers processing video clips) inherently rely on temporal data. For instance, methods explaining predictions based on temporal attention weights or gradients through time cannot be applied to single images.
Reasoning-based methods often build representations like scene graphs or causal graphs that explicitly model temporal relationships and interactions between objects over time [87, 88, 89, 90]. While scene graphs can be constructed for static images, they lack the temporal dimension crucial for explaining event-based anomalies (e.g., explaining why a sequence of actions is abnormal). The richness of context required for meaningful reasoning-based explanations is often more readily available in video sequences than in isolated images. For example, [89] constructs a scene graph capturing global interactions and uses metric learning on (subject, predicate, object) triplets derived from video to identify anomalous interactions, providing interpretable feedback on which interaction deviates from learned norms.
Methods utilizing optical flow [6] explicitly model motion. Explanations derived from optical flow analysis (e.g., identifying regions with anomalous motion patterns) are inherently video-specific.
Intrinsic interpretable methods using attributes like velocity, pose trajectory, or specific action labels [96, 97, 99, 98] rely on temporal information to define and measure these attributes. While some attributes (like object class) exist in images, many key attributes for explaining video anomalies (e.g., running speed, interaction type) do not.
In essence, explainable VAD methods often exploit the temporal dimension and dynamic content unique to videos, making their direct application or the relevance of their explanations limited in the context of static images.
IV-C Commonalities and Transferable Approaches
Despite the differences, certain explainability paradigms can bridge the gap between IAD and VAD, often requiring adaptation.